OpenStudy (anonymous):

A 32 ounce bottle of water costs $1.60. What is the cost per ounce?

OpenStudy (anonymous):

.05 CENTS PER ONCE

OpenStudy (anonymous):

160/32

OpenStudy (tkhunny):

The very words "something per something else" should tell you we have a division problem. \(Cost\;per\;Ounce = \dfrac{Some\;Cost\;Measure}{Some\;Ounce\;Measure} = \dfrac{$1.60}{32\;oz} = \dfrac{1.60}{32}\;$/oz = \dfrac{160}{32}\;cent/oz\)